please help me solve this exponents sum
x^x^x^x^infinite=4
if x to the power x to the power x to the power upto infinity=4 find x

Answer by sachi(548)

x^x^x^x^infinite=4------1
or x^(x^x^x^infinite)=4
or x^4=4 by eqn 1 above
taking log on both sides
4logx=log4
or logx=(1/4)log4=log4^(1/4)
or x=4^(1/4) ans